factoring and solving
4a to the 2 power-9=0

Answer by stanbon(75887)

4a^2-9=0
This form is called "the difference of squares".
It factors as follows: (2a-3)(2a+3)=0
a=3/2 or a=-3/2
Cheers,
Stan H.